Please help! Don't have time. Due 11:59pm 02/29. Thank you!

Required information The Foundational 15 (LO4-1, LO4-2, LO4-3, LO4-4, LO4-5) [The following information applies to the questi11. What is the cost of ending work in process inventory for conversion? (Round your intermediate calculations to 2 places.)12. What is the cost of materials transferred to finished goods? (Round your intermediate calculations to 2 places.) Cost of13. What is the amount of conversion cost transferred to finished goods? (Round your intermediate calculations to 2 places.)14. Prepare the journal entry to record the transfer of costs from Work in Process to Finished Goods. (If no entry is requireFoundational 4-15 15-a. What is the total cost to be accounted for? 15-b. What is the total cost accounted for? Total cost to
Answer #1

Unit completed and transferred out = 5100+37600-8200 = 34500

Equivalent unit of material = 34500+8200*100% = 42700

Equivalent unit of conversion = 34500+8200*50% = 38600

Cost per equivalent unit of material = (18570+120205)/42700 = 3.25

Cost per equivalent unit of conversion = (16430+80500+98000)/38600 = 5.05

11) Cost of ending WIP for conversion = 4600*5.05 = 23230

12) Cost material transferred to finished goods = 34500*3.25 = 112125

13) Cost of conversion transferred to finished goods = 34500*5.05 = 174225

14) Journal entry

Date General Journal Debit Credit
Finished goods (112125+174225) 286350
Work in process 286350

15a) Total Cost to be accounted for = 35000+298705 = 333705

15b) Total Cost accounted for = 286350+47355 = 333705